OpenStudy (anonymous):

More words.. A wheel on Elsa's bicycle is .9m in diameter. Elsa races the bicycle for 79m . How many times does the wheel turn as the bicycle travels this distance? Use the value 3.14 for pi Round your answer to the nearest tenth. Do not round any intermediate steps.

OpenStudy (anonymous):

How would I set up the equation?

OpenStudy (anonymous):

find circumference. divide 79 by this answer.

OpenStudy (anonymous):

for every round completed by the wheel, a distance of the wheel's circumference is traversed, which is 2pi*r. diameter=0.9 radius =0.45 79=2*3.14*.45*no.of revolutions no.of revolutions= 79/(2*.45*3.14) =27.95

OpenStudy (anonymous):

Ugh.. that seems so easy.. Wish I wasn't so stupid when words are thrown at me >.< Thank you guys so much!

OpenStudy (anonymous):

Ermm one more thing.. I have to round to the nearest tenth.. I"m not sure how this should be done because it seems if I did that, the answer would be 28 but it wouldn't complete 28 revolutions so the answer would be wrong.. Should I just leave it 27.9?

OpenStudy (anonymous):

I think .95 would get round to the next higher number, so in this case the answer should be 28. In general i'd say .5 would be rounded off to 1

OpenStudy (anonymous):

Ok, I will do just that.. Thank you again good sir

OpenStudy (anonymous):

Glad to help! :)